OpenStudy (anonymous):

How would you get the points and graph for this equation? y=2/3x?

OpenStudy (katrinakaif):

y = mx + b y = 2/3x + 0 (as you don't have a 'b' B = Y-intercept Start at (0,0) - Make a point Since your slope is 2/3 (rise/run) move up 2 units, and 3 units to the right. Make another point From there, move 2 up again, and 3 to the right..Keep repeating the process till you have a long enough line. =)

OpenStudy (anonymous):

thank youu!(:

OpenStudy (katrinakaif):

Anytime =D